According to a polling​ organization,

2424​%

of adults in a large region consider themselves to be liberal. A survey asked 200 respondents to disclose their political​ philosophy: Conservative,​ Liberal, Moderate. Treat the results of the survey as a random sample of adults in this region. Do the survey results suggest the proportion is higher than that reported by the polling​ organization? Use an

alphaαequals=0.050.05

level of significance.

Because np 0 left parenthesis 1 minus p 0 right np01−p0) equals= blank

blank ​10; the sample size is blank

less than

greater than

​5% of the population​ size, and the sample

can be reasonably assumed to be random,

cannot be reasonably assumed to be random,

is given to be random,

is given to not be random,

the requirements for testing the hypothesis

are

are not

satisfied.

​(Round to one decimal place as​ needed.)
Answer #1

np(1 - p) = 200 * 0.24 * (1 - 0.24) = 36.48

Hence,

The correct answer in blanks will be:

36.48; Greater than; Less than 5%; Can be reasonably assumed to be normal; are